# Fusion Systems related to polynomial representations of $\operatorname{SL}_{2}(q)$

### v1

@@ 4. Polynomial fusion systems
Let $p$ be an odd prime, $11\leqslant n\leqslant p-$ 1, $S=S_{n}(q)$ and define subgroups $R=Z(S)U$ and $Q=Z_{2}(S)U$. Then $R$ is elementary abelian of order $q^{2}$ and $Q\cong S_{1}(q)$ by Corollary 2.9. In [CP10, Section 4] two groups $P_{R}$ and $P_{Q}$ are defined. These groups have a normal subgroup isomorphic to $R$ and $Q$ respectively with $\mathrm{Aut}_{P_{R}}(R)\cong\mathrm{GL}_{2}(q)\cong\mathrm{Out}_{P_{Q}}(Q)$. For a moment set $P=D\ltimes V_{n}(q)$. Then monomorphisms $\Psi_{R}:N_{P}(R)\rightarrow P_{R}$ and $\Psi_{Q}:N_{P}(Q)\rightarrow P_{Q}$ are given such that the free amalgamated products $F(n,\mathbb{K},R)=P\ast_{N_{P}(R)}P_{R}$ and $F(n,\mathbb{K},Q)=P\ast_{N_{P}(Q)}P_{Q}$ can be constructed. Then $P$ and $P_{R}$, respectively $P_{Q}$, are identified as subgroups of these products. In particular, $S$ is a subgroup and the fusion systems $\mathcal{F}(n,q,R)=\mathcal{F}_{S}(F(n,\mathbb{K},R))$ and $\mathcal{F}(n,q,Q)=\mathcal{F}_{S}(F(n,\mathbb{K},Q))$ are shown to be saturated and mostly exotic.
Now write $P^{*}=D^{*}\rtimes V_{n}(q)$, and note that since any element of $\mathrm{Aut}(\mathbb{K})$ acts on $P_{R}$, $P_{Q}$ by raising matrix entries to some appropriate $p$-power, the maps $\Psi_{R}$ and $\Psi_{Q}$ defined above are $\mathrm{Aut}(\mathbb{K})$-equivariant. Thus, in parallel to the construction in Section 3, one may similarly define
@@
and maps $\Psi_{Q}^{*}:N_{P^{*}}(Q)\rightarrow P_{Q}^{*}$ and $\Psi_{R}^{*}:N_{P^{*}}(R)\rightarrow P_{R}^{*}$ such that the free amalgamated products
$F^{*}(n,\mathbb{K},R):=P*_{N_{P^{*}}(R)}P^{*}_{R}\mbox{ and }F(n,\mathbb{K},Q):=P*_{N_{P^{*}}(Q)}P^{*}_{Q}$
can be constructed. We write $\mathcal{F}^{*}(n,q,Q)$, respectively $\mathcal{F}^{*}(n,q,R)$, for the associated fusion systems, regarded as the largest saturated fusion systems defined on $S_{n}(q)$, where $q=p^{m}>p$, among those such that ()=$\mathcal{E}(\mathcal{F})=\{V\}\cup Q^{S}$, respectively ()=$\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$. It is clear that $\mathcal{F}(n,q,X)$ is a fusion subsystem of $\mathcal{F}^{*}(n,q,X)$ for $X\in\{R,Q\}$.
@@
Note that neither the isomorphism type of $S$ nor the set of $\mathcal{F}_{S}(S)$-centric subgroups is affected by appending $p^{\prime}$-automorphisms to automisers of subgroups in a fusion system. Furthermore, the essential subgroups remain unchanged when passing from $\mathcal{F}_{S}(P)$ to $\mathcal{F}_{S}(P^{*})$, and from $\mathcal{F}_{N_{S}(X)}(P_{X})$ to $\mathcal{F}_{N_{S}(X)}(P_{X}^{*})$ where $X\in\{R,Q\}$. Therefore, the proof of [CP10, Theorem 4.9] applies to this situation too. Since $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,X)}(X))=O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}(n,q,X)}(X))$ for $X\in\{V,R,Q\}$, it quickly follows from [AKO11, Lemma I.7.6] that $\mathcal{F}(n,q,X)$ has $p^{\prime}$-index in $\mathcal{F}^{*}(n,q,X)$ for $X\in\{R,Q\}$. The remaining points follow easily. ∎
@@
Now observe that by [PS21, Lemma 6.4] the fusion systems $\mathcal{F}^{*}(n,q,Q)$ and $\mathcal{F}^{*}(n,q,R)$ determine further systems via a process known as “pruning”. More precisely, a subset of the automorphisms of the essential subgroup $V$ can be removed in either case to form, respectively, two new saturated fusion systems which we call $\mathcal{F}^{*}(n,q,Q)_{P}$ and $\mathcal{F}^{*}(n,q,R)_{P}$ on $S_{n}(q)$ with
@@
(note that closely related examples were previously considered in [HLL23, Lemma 5.10]). Analogously we write $\mathcal{F}_{\Lambda}^{*}(q)_{P}$ for the system obtained from $\mathcal{F}_{\Lambda}^{*}(q)$ by pruning $\Lambda(q)$ so that
@@
Suppose that $11<n\leqslant p-$ 1 and $\mathcal{F}=\mathcal{F}^{*}(n,q,R)_{P}$. Then the following hold:
@@
    $O^{p^{\prime}}(\mathcal{F})$ is simple unless $n=p-2$ and $q=p>3$, in which case
@@
We observe that $R\in\mathcal{E}(\mathcal{F}^{*}(n,q,R)_{P})$ and as $n>1$, $R\not\trianglelefteq S$. Since $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}(R))\cong\operatorname{SL}_{2}(q)$ acts irreducibly on $R$, [AKO11, Proposition I.4.5] implies that $O_{p}(\mathcal{F})=1$. We observe that ${\mathcal{F}^{*}(n,q,R)_{P}}^{frc}=\{S\}\cup R^{S}$ and so, in order to calculate $\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)_{P}}^{0}(S)$ via [AKO11, Theorem I.7.7], it suffices to lift elements only from $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)_{P}}(R))\cong\operatorname{SL}_{2}(q)$. Hence, $|\mathrm{Out}_{\mathcal{F}^{*}(n,q,R)_{P}}^{0}(S)|=q-1$.
@@
Let $X$ be a proper non-trivial strongly closed subgroup of $\mathcal{F}^{*}(n,q,R)_{P}$. Then $X\trianglelefteq S$ so that $1\neq X\cap Z(S)$ so that $1\neq X\cap R$. Again, as $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}(R))$ acts irreducibly on $R$, we deduce that $R\leq X$ so that $X[V,S]=\langle X^{S}\rangle R$. Let $k$ be an element of $\mathrm{Aut}^{0}_{\mathcal{F}^{*}(n,q,R)_{P}}(S)$ of order $q-1$. Then we may arrange that $k$ is the lift of an element of $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)_{P}}(R))}(\mathrm{Aut}_{S}(R))$ and so acts as $\lambda^{-1}$ on $S/V$ and as $\lambda$ on $Z(S)$ for some $\lambda\in\mathbb{K}^{*}$. Using the injectivity of the map $\delta$ when restricted to $p^{\prime}$-elements from Lemma 2.20, such an element acts in the same manner as an element of $\mathrm{Aut}_{P^{*}}(S)$ which may be represented in the form $\left(1,\lambda,\left(\begin{matrix}1&0\\ 0&\lambda\end{matrix}\right)\right)$. Hence, $k$ acts on $V/[V,S]$ as $\lambda^{n+1}$, so that $k$ is irreducible on $V/[V,S]$. Thus, we deduce that $X=R[V,S]$ is the unique proper non-trivial strongly closed subgroup of $\mathcal{F}^{*}(n,q,R)_{P}$.
@@
Now if $\mathcal{N}\trianglelefteq O^{p^{\prime}}(\mathcal{F}^{*}(n,q,R)_{P})$ is supported on $R[V,S]$ then $\mathrm{Aut}_{\mathcal{N}}(R[V,S])\trianglelefteq\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)_{P}}(R[V,S])$. It follows that $\mathrm{Out}_{S}(R[V,S])$ must commute with $\langle k|_{R[V,S]}\rangle\mathrm{Inn}(R[V,S])/\mathrm{Inn}(R[V,S])$, from which we deduce that $k$ centralises $V/[V,S]$. This occurs only if $q=p>3$ and $n=p-2$ and in this case, $\mathcal{N}=O^{p}(O^{p^{\prime}}(\mathcal{F}^{*}(p-2,p,R)_{P}))\cong O^{p^{\prime}}(\mathcal{F}^{*}(p-3,p,R)_{P})$ using that $R[V,S]\cong S_{n-1}(q)$ by Lemma 2.8, and then applying Theorem 1.1. In every other instance, we have that $O^{p^{\prime}}(\mathcal{F}^{*}(n,q,R)_{P})$ is simple and no normal subsystem of $\mathcal{F}^{*}(n,q,R)_{P}$ is supported on $R[V,S]$. Applying [vB22, Theorem 3.12], we see that $O^{p^{\prime}}(\mathcal{F}^{*}(n,q,R)_{P})$ is exotic unless perhaps $q=p>3$ and $n=p-2$. But in this case, if $O^{p^{\prime}}(\mathcal{F}^{*}(p-2,p,R)_{P})$ was realised by $G$ then $O^{p}(O^{p^{\prime}}(\mathcal{F}^{*}(p-2,p,R)_{P}))$ is a saturated fusion system supported on $O^{p}(G)\cap S\in\mathrm{Syl}_{p}(O^{p}(G))$ by the hyperfocal subgroup theorem of Puig [Pui00, §1.1]. By Theorem 1.1, $\mathcal{F}_{O^{p}(G)\cap S}(O^{p}(G))\cong O^{p^{\prime}}(\mathcal{F}^{*}(p-3,p,R)_{P})$ is exotic, an obvious contradiction. Hence, $O^{p^{\prime}}(\mathcal{F}^{*}(p-2,p,R)_{P})$ is also exotic. ∎
The fusion systems introduced in this section, together with those introduced in Section 3 are collectively called polynomial fusion systems.
@@
    $O_{p}(\mathcal{F}^{*}(n,q,Q)_{P})=Z(S)$ and $\mathcal{F}^{*}(n,q,Q)_{P}/Z(S)\cong\mathcal{F}^{*}(n-1,q,R)_{P}$.
@@
To prove that $\mathcal{N}_{i}$ is saturated we adopt the same method as in the proof of [Gra18, Lemma 3.7], which ultimately relies on [BLO06, Theorem 5.1]. The equlalites $O_{p}(\mathcal{N}^{i})=1$ and $\mathcal{E}(\mathcal{N}^{i})=\{R^{N^{i}}\}$ are both clear by construction and applying [AKO11, Proposition I.4.5]. All that remains is to show that $\mathcal{N}^{i}\cong\mathcal{F}^{*}(i,q,R)_{P}$, but this is immediate from the fact that $N^{i}\cong S_{i}(q)$ (by Lemma 2.11(6)) and our main Theorem 1.1. ∎

### v2

@@ 4. Polynomial fusion systems
Let $p$ be an odd prime and for $1\leqslant n\leqslant p-1$ set $S=S_{n}(q)$. Recall the definition of $U$ from Notation 2.2 and the definition of $R$ and $Q$ from Notation 2.21. Thus $R$ is elementary abelian of order $q^{2}$ and $Q\cong S_{1}(q)$ by Corollary 2.13. In [CP10, Section 4] two groups $P_{R}$ and $P_{Q}$ are defined. These groups have a normal subgroup isomorphic to $R$ and $Q$ respectively with $\mathrm{Aut}_{P_{R}}(R)\cong\mathrm{GL}_{2}(q)\cong\mathrm{Out}_{P_{Q}}(Q)$. Recall from Notation 2.3 that $P_{n}(q)=D\ltimes V_{n}(q)$. Then monomorphisms $\Psi_{R}:N_{P_{n}(q)}(R)\rightarrow P_{R}$ and $\Psi_{Q}:N_{P_{n}(q)}(Q)\rightarrow P_{Q}$ are given such that the free amalgamated products $F(n,\mathbb{K},R)=P_{n}(q)\ast_{N_{P_{n}(q)}(R)}P_{R}$ and $F(n,\mathbb{K},Q)=P_{n}(q)\ast_{N_{P_{n}(q)}(Q)}P_{Q}$ can be constructed. Then $P_{n}(q)$ and $P_{R}$, respectively $P_{Q}$, are identified as subgroups of these products. In particular, $S$ is a subgroup and the fusion systems $\mathcal{F}(n,q,R)=\mathcal{F}_{S}(F(n,\mathbb{K},R))$ and $\mathcal{F}(n,q,Q)=\mathcal{F}_{S}(F(n,\mathbb{K},Q))$ are shown to be saturated and mostly exotic.
Recall from Notation 2.3 that $P_{n}^{*}(q)=D^{*}\rtimes V_{n}(q)$, and note that since any element of $\mathrm{Aut}(\mathbb{K})$ acts on $P_{R}$, $P_{Q}$ by raising matrix entries to some appropriate $p$-power, the maps $\Psi_{R}$ and $\Psi_{Q}$ defined above are $\mathrm{Aut}(\mathbb{K})$-equivariant. Thus, in parallel to the construction in Section 3, one may similarly define
@@
and maps $\Psi_{Q}^{*}:N_{P_{n}^{*}(q)}(Q)\rightarrow P_{Q}^{*}$ and $\Psi_{R}^{*}:N_{P^{*}_{n}(q)}(R)\rightarrow P_{R}^{*}$ such that the free amalgamated products
$F^{*}(n,\mathbb{K},R):=P_{n}^{*}(q)*_{N_{P_{n}^{*}(q)}(R)}P^{*}_{R}\mbox{ and }F^{*}(n,\mathbb{K},Q):=P_{n}^{*}(q)*_{N_{P_{n}^{*}(q)}(Q)}P^{*}_{Q}$
can be constructed.
Notation 4.1.
We write $\mathcal{F}^{*}(n,q,R)$, respectively $\mathcal{F}^{*}(n,q,Q)$, for the fusion systems associated to $F^{*}(n,\mathbb{K},R)$, respectively $F^{*}(n,\mathbb{K},Q)$, which may be regarded as the largest saturated fusion systems defined on $S_{n}(q)$, where $q=p^{m}>p$, among those such that $\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$, respectively $\mathcal{E}(\mathcal{F})=\{V\}\cup Q^{S}$.
It is clear that $\mathcal{F}(n,q,X)$ is a fusion subsystem of $\mathcal{F}^{*}(n,q,X)$ for $X\in\{R,Q\}$.
@@
Note that neither the isomorphism type of $S$ nor the set of $\mathcal{F}_{S}(S)$-centric subgroups is affected by appending $p^{\prime}$-automorphisms to automisers of subgroups in a fusion system. Furthermore, the essential subgroups remain unchanged when passing from $\mathcal{F}_{S}(P_{n}(q))$ to $\mathcal{F}_{S}(P^{*}_{n}(q))$, and from $\mathcal{F}_{N_{S}(X)}(P_{X})$ to $\mathcal{F}_{N_{S}(X)}(P_{X}^{*})$ where $X\in\{R,Q\}$. Therefore, the proof of [CP10, Theorem 4.9] applies to this situation too. Since $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,X)}(X))=O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}(n,q,X)}(X))$ for $X\in\{V,R,Q\}$, it quickly follows from [AKO11, Lemma I.7.6] that $\mathcal{F}(n,q,X)$ has $p^{\prime}$-index in $\mathcal{F}^{*}(n,q,X)$ for $X\in\{R,Q\}$. The remaining points follow easily. ∎
@@
Now observe that by [PS21, Lemma 6.4] the fusion systems $\mathcal{F}^{*}(n,q,Q)$ and $\mathcal{F}^{*}(n,q,R)$ determine further systems via a process known as “pruning”. More precisely, a subset of the automorphisms of the essential subgroup $V$ can be removed in either case to form, respectively, two new saturated fusion systems.
Notation 4.4.
Define the saturated fusion systems $\mathcal{F}^{*}(n,q,Q)_{P}$ and $\mathcal{F}^{*}(n,q,R)_{P}$ on $S_{n}(q)$ to be those obtained by pruning $V$ from $\mathcal{F}^{*}(n,q,Q)$ and $\mathcal{F}^{*}(n,q,R)$. Then
@@
(note that closely related examples were previously considered in [HLL23, Lemma 5.10]).
Analogously we write $\mathcal{F}_{\Lambda}^{*}(q)_{P}$ for the system obtained from $\mathcal{F}_{\Lambda}^{*}(q)$ by pruning $\Lambda(q)$ (see Notation 3.3) so that
@@
Suppose that $1<n\leqslant p-1$ and $\mathcal{F}=\mathcal{F}^{*}(n,q,R)_{P}$. Then the following hold:
@@
    $O^{p^{\prime}}(\mathcal{F})$ is simple unless n=p2n=p-2 and $q=p>3$, in which case
@@
We observe that $R\in\mathcal{E}(\mathcal{F}^{*}(n,q,R)_{P})$ and, as $n>1$, we see that $R\not\trianglelefteq S$. Since $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}(R))\cong\operatorname{SL}_{2}(q)$ acts irreducibly on $R$, [AKO11, Proposition I.4.5] implies that $O_{p}(\mathcal{F})=1$. We observe that ${\mathcal{F}^{*}(n,q,R)_{P}}^{frc}=\{S\}\cup R^{S}$ and so, in order to calculate $\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)_{P}}^{0}(S)$ via [AKO11, Theorem I.7.7], it suffices to lift elements only from $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)_{P}}(R))\cong\operatorname{SL}_{2}(q)$. Hence, $|\mathrm{Out}_{\mathcal{F}^{*}(n,q,R)_{P}}^{0}(S)|=q-1$.
@@
Let $X$ be a proper non-trivial strongly closed subgroup of $\mathcal{F}^{*}(n,q,R)_{P}$. Then $X\trianglelefteq S$ so that $1\neq X\cap Z(S)$ so that $1\neq X\cap R$. Again, as $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}(R))$ acts irreducibly on $R$, we deduce that $R\leq X$ so that $X[V,S]=\langle X^{S}\rangle R$. Let $k$ be an element of $\mathrm{Aut}^{0}_{\mathcal{F}^{*}(n,q,R)_{P}}(S)$ of order q1q-1. Then we may arrange that $k$ is the lift of an element of $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)_{P}}(R))}(\mathrm{Aut}_{S}(R))$ and so acts as $\lambda^{-1}$ on $S/V$ and as $\lambda$ on $Z(S)$ for some $\lambda\in\mathbb{K}^{*}$. Recall the map $\delta$ from Notation 2.25. Using the injectivity of $\delta$ when restricted to $p^{\prime}$-elements from Lemma 2.26, such an element acts in the same manner as an element of $\mathrm{Aut}_{P^{*}_{n}(q)}(S)$ which may be represented in the form $\left(1,\lambda,\left(\begin{matrix}1&0\\ 0&\lambda\end{matrix}\right)\right)$. Hence, $k$ acts on $V/[V,S]$ as $\lambda^{n+1}$, so that $k$ is irreducible on $V/[V,S]$. Thus, we deduce that $X=R[V,S]$ is the unique proper non-trivial strongly closed subgroup of $\mathcal{F}^{*}(n,q,R)_{P}$.
@@
Now if $\mathcal{N}\trianglelefteq O^{p^{\prime}}(\mathcal{F}^{*}(n,q,R)_{P})$ is supported on $R[V,S]$ then $\mathrm{Aut}_{\mathcal{N}}(R[V,S])\trianglelefteq\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)_{P}}(R[V,S])$. It follows that $\mathrm{Out}_{S}(R[V,S])$ must commute with $\langle k|_{R[V,S]}\rangle\mathrm{Inn}(R[V,S])/\mathrm{Inn}(R[V,S])$, from which we deduce that $k$ centralises $V/[V,S]$. This occurs only if $q=p>3$ and n=p2n=p-2 and in this case, $\mathcal{N}=O^{p}(O^{p^{\prime}}(\mathcal{F}^{*}(p-2,p,R)_{P}))\cong O^{p^{\prime}}(\mathcal{F}^{*}(p-3,p,R)_{P})$ using that $R[V,S]\cong S_{n-1}(q)$ by Lemma 2.12, and then applying Theorem 1.1. In every other instance, we have that $O^{p^{\prime}}(\mathcal{F}^{*}(n,q,R)_{P})$ is simple and no normal subsystem of $\mathcal{F}^{*}(n,q,R)_{P}$ is supported on $R[V,S]$. Applying [VAN25, Theorem 3.12], we see that $O^{p^{\prime}}(\mathcal{F}^{*}(n,q,R)_{P})$ is exotic unless perhaps $q=p>3$ and n=p2n=p-2. But in this case, if $O^{p^{\prime}}(\mathcal{F}^{*}(p-2,p,R)_{P})$ was realised by $G$ then $O^{p}(O^{p^{\prime}}(\mathcal{F}^{*}(p-2,p,R)_{P}))$ is a saturated fusion system supported on $O^{p}(G)\cap S\in\mathrm{Syl}_{p}(O^{p}(G))$ by the hyperfocal subgroup theorem of Puig [PUI00, §1.1]. By Theorem 1.1, $\mathcal{F}_{O^{p}(G)\cap S}(O^{p}(G))\cong O^{p^{\prime}}(\mathcal{F}^{*}(p-3,p,R)_{P})$ is exotic, an obvious contradiction. Hence, $O^{p^{\prime}}(\mathcal{F}^{*}(p-2,p,R)_{P})$ is also exotic. ∎
Definition 4.6.
A saturated fusion system $\mathcal{F}$ is a polynomial fusion system if $\mathcal{F}$ is a core-free subsystem of $p^{\prime}$-index in any of $\mathcal{F}^{*}(n,q,R)$, $\mathcal{F}^{*}(n,q,Q)$, $\mathcal{F}^{*}(n,q,R)_{P}$ or $\mathcal{F}_{\Lambda}^{*}(q)$, where $1\leq n<p$ and $q>p$.
We note that $\mathcal{F}^{*}(1,q,R)\cong\mathcal{F}_{S}(G)$ where $G\cong O^{p}(\mathrm{P}\Gamma\mathrm{L}_{3}(q))$, $O_{p}(\mathcal{F}^{*}(1,q,R)_{P})\neq 1$, and $\mathcal{F}^{*}(2,q,Q)\cong\mathcal{F}_{S}(G)$ where $G\cong O^{p}(\mathrm{Aut}(\mathrm{PSp}_{4}(q)))$.
@@
    $O_{p}(\mathcal{F}^{*}(n,q,Q)_{P})=Z(S_{n}(q))$ and $\mathcal{F}^{*}(n,q,Q)_{P}/Z(S_{n}(q))\cong\mathcal{F}^{*}(n-1,q,R)_{P}$.
@@
To prove that $\mathcal{N}_{i}$ is saturated we adopt the same method as in the proof of [GRA18, Lemma 3.7], which ultimately relies on [BLO06, Theorem 5.1]. The equalities $O_{p}(\mathcal{N}^{i})=1$ and $\mathcal{E}(\mathcal{N}^{i})=\{R^{N^{i}}\}$ are both clear by construction and applying [AKO11, Proposition I.4.5]. All that remains is to show that $\mathcal{N}^{i}\cong\mathcal{F}^{*}(i,q,R)_{P}$, but this is immediate from the fact that $N^{i}\cong S_{i}(q)$ (by Lemma 2.15(6)) and our main Theorem 1.1. ∎